VEGAS Team Meeting

Attendees: Amanda, Anish, Paul, Joe, Marty, Richard, Mark, Randy, Dan, John, Hong

Agenda:

  1. OGP calibration; status of ADC calibration in general
  2. Usage of MMCM in low bandwidth mode
  3. ADC calibration routines and documentation for DIBAS
  4. Mode 10-19 (ghost frequencies issue)
  5. Blanking of last switched integration
  6. Status of “Manager writes to many sub-bands in 8 subband mode” problem (release and test)
  7. Content of August 'release'
  8. Use of test time (including overnights)
  9. Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues
  10. Resource plan for Q4 FY2015

Meeting notes:

  1. OGP calibration; status of ADC calibration in general
    • we are basically done
    • the MR has been signed by the sponsor (Jason Ray)
      • Not an effect on August release
  2. Usage of MMCM in low bandwidth mode: VEGAS, unfortunately, has some modes which clock the ADC at very low sample rates, which forces us to use low bandwidth mode
    • there is still the issue about this fundamental flaw with the MMCM calibrations occasionally failing: we discussed at the end of the last meeting ways we could get around this, but I haven't heard any more details on it and I don't have any action items to follow through on. Right now our Manager expects that these measurements will occasionally fail.
  3. ADC calibration routines and documentation for DIBAS
    • we are putting the finishing touches on the docs intended for DIBAS
  4. Mode 10-19 (ghost frequencies issue)
    • Randy looking at bof files
    • Joe on GPU with time series snapshots - working on issue with extraction routine
      • real and ghost spectra are sometimes swapped
    • Anish looking at data from above
  5. Blanking of last switched integration
    • Released
  6. Status of “Manager writes to many sub-bands in 8 subband mode” problem (release and test)
    • Released
  7. Use of test time (including overnights)
    • Unscheduled request(s)
      • none
    • Tuesday
      • 13:00 - 15:00 mixed mode testing and checking vegas exposure time with LO1 switching [Roshi]
    • Wednesday
      • No time yet reserved
    • Thursday
      • 8:00am – 12pm; Streaming Data Display tests [Kepley/Brandt]
  8. Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues

Other

  1. Please make sure to properly code your timesheets for the current time period and going forward.
    • Software engineers should be using a GBT ops software CM&E account - Account 893252
    • Electronic engineers should be using a GBT ops electronics CM&E account - Account 432510
    • Commissioning activities should be getting charged to hardware support and testing - Account 432510
    • Scientists assisting with shared-risk commissioning should be charging an observer support account
      • 432540 for the most part. Accounts 142020 (GB-based staff) and 142010 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)

-- MartyBloss - 2014-06-23
Topic revision: r2 - 2014-06-23, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback